OpenStudy (anonymous):

solve: lnx^3-2(lnx)=8

OpenStudy (anonymous):

lnx^3-2(lnx)=8 3lnx-2(lnx)=8 1ln x =8 ln x=8 take e power on both sides we get e^ln x=e^8 x=e^8 use calculator..to get x=2980.957987 ans....

OpenStudy (anonymous):

any question missjj

OpenStudy (anonymous):

think i got it. TY

OpenStudy (anonymous):

ok..yw